OpenStudy (anonymous):

how do I take the implicit derivative of 4cosXsinY=1 ?

OpenStudy (anonymous):

you'd have to take the d/dx of each side so you're going to have to use the product rule

OpenStudy (anonymous):

4cos(x) = u sin(y)= v

OpenStudy (anonymous):

u'= -4sin(x)

OpenStudy (anonymous):

Right, I got down to where I have 4(sinx-cos(dy/dx)) + (cosx)(-cos(dy/dx)) + (sinx)(siny) and now I don't really know if that's right

OpenStudy (anonymous):

Do you have any idea if or where I messed up?

OpenStudy (anonymous):

what did you get before that

OpenStudy (anonymous):

what did you get for the derivative of cos(y)

OpenStudy (anonymous):

the answer is supposed to be Y' = tanXtanY

OpenStudy (anonymous):

The first thing I did was to take the the derivitive of 4cosXsinY=1 so that the first thing I did was to come up with 4(sinx-cos(dy/dx)) + [(cosx)(siny)' + (cosx)'(siny)]

OpenStudy (anonymous):

i think that is where you got off track d/dx[sin(y)]=cos(y)dy/dx

OpenStudy (anonymous):

Then the second thing I did was the equation I gave earlier: 4(sinx-cos(dy/dx)) + (cosx)(-cos(dy/dx)) + (sinx)(siny)

OpenStudy (anonymous):

look at cos(y) as cos(u) i guess you could say so the derivative of this would be cos(u)u' u=y u'=dy/dx

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

i mean't sin(y) derivative = cos(u)u'

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

It has been a while since I had trig too so that isn't really helping too much either

OpenStudy (anonymous):

so you should get 4cos(x)(cos(y)y-4sin(y)sin(x)=0 *add the 4sin(y)sin(x) to the left and factor out y' on the left y'(4cos(x)cos(y))=4sin(y)sin(x) *divide by 4cos(x)cos(y) y'=tanYtanX

OpenStudy (anonymous):

or TanXTanY same thing

OpenStudy (anonymous):

Thanks I really appreciate it! I'm brand new to this and I think I should change my name, do you know how to do that too?

OpenStudy (anonymous):

have no idea You might have to send them something or what not

OpenStudy (anonymous):

Thanks, I need to get back to studying now, again I thank you for the help